OpenStudy (anonymous):

find cos (2sin^-1(5/7))

OpenStudy (kc_kennylau):

Draw a triangle representing \(\sin^{-1}\dfrac57\) first

OpenStudy (anonymous):

is it possible for someone to help break this problem down for me? i need to find the exact values. thanks in advance!

OpenStudy (anonymous):

Kc_kennylau.. How do I draw a triangle with the sin^-1? does that come out to 1/sin?

OpenStudy (kc_kennylau):

First you have to know what \(\sin^{-1}\) means.

OpenStudy (kc_kennylau):

It does not mean \(\dfrac1\sin\), but instead it's the inverse function of \(\sin\)

OpenStudy (kc_kennylau):

For example, if \(\sin30^\circ=0.5\), then \(\sin^{-1}0.5=30^\circ\)

OpenStudy (kc_kennylau):

Draw an right triangle where \(\sin\theta=\dfrac57\) first :)

OpenStudy (anonymous):

awesome, going to do so now!!

OpenStudy (kc_kennylau):

Okay, and as always you can use the draw button below :)

OpenStudy (anonymous):